Molecular vibrations

Ask Hjorth Larsen

Center for Atomic-scale Materials Design

Molecular vibrations

dfjk

Metamaterials

Background

▶ Metamaterial: properties determined by large-scale structure,

not chemical composition ▶ Structures can affect light if scale is small compared to wave

length ▶ We can make all sorts of weird structures

Optical interactions

▶ Generally, refractive index is n = \sqrt{ϵμ} ▶ μ, ϵ > 0: the usual case ▶ μ < 0 xor ϵ < 0: refractive index is complex, amplitude decays

with propagation in material ▶ μ, ϵ < 0: strategic abuse of power laws permits n < 0